OpenStudy (anonymous):

Suppose f is a differentiable function such that f'(x)<=3 for all x E(2,7). if f(-2)=3 the Mean Value Theorem says that f(7)<=V for what value of V? (Choose V as small as possible)

OpenStudy (anonymous):

is it \((2,7)\) or \((-2,7)\) ?

OpenStudy (anonymous):

you have as an interval \((2,7)\) but then you have \(f(-2)\) so there is some confusion here

OpenStudy (anonymous):

its -2 sorry lol

OpenStudy (anonymous):

ok how far from \(-2\) to \(7\) ?

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):



OpenStudy (anonymous):

9

OpenStudy (anonymous):

whew ok so imagine you are walking, but you cannot walk faster than 3 mph you walk for 9 hours what is the furthest you can go?

OpenStudy (anonymous):

27

OpenStudy (anonymous):

right and you start at mile maker \(3\) so if you go at most 27 miles, what mile marker could you reach?

OpenStudy (anonymous):

would it be 30 ?

OpenStudy (anonymous):

yes and that is one way to interpret the problem, and probably do it in your head we can do it the other way if you like

OpenStudy (anonymous):

is there an easier way lol

OpenStudy (anonymous):

how did u know i was walking 3 mph

OpenStudy (anonymous):

and how did u know marker was set at 3

OpenStudy (anonymous):

\[f'(x)<3\] is like saying your speed is at most 3 mph \(f(-2)=3\) means when you start counting you are at mile marker 3, and since you cannot walk more that 27 miles in 9 hours, you can be at most at mile marker 30 now lets do it the calc way

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

by the mvthm \[\frac{f(7)-f(-2)}{7-(-2)}=f'(c)\] for some \(c\) but since \(f'(x)<3\) we have \[\frac{f(7)-f(-2)}{7-(-2)}\leq 3\]

OpenStudy (anonymous):

you know \(f(-2)=3\) and so this is the same as \[\frac{f(7)-3)}{9}\leq 3\] or \[f(7)-3\leq 27\] or \[f(7)\leq 30\]

OpenStudy (anonymous):

oh how cool this is shorter in a way

OpenStudy (anonymous):

is this Rolls formula?